# Fast and low energy approximate full adder based on FELIX logicthanks: Preprint Submitted to Journal of Engineering

### v2

@@ 4.1 Circuit-Level Simulation And Analysis
As shown in Table 8, the memristors’ count in both approximate scenarios and implementation approaches equals the memristors’ count in the exact 8-bit structure [16]. In both scenarios, the approximate 8-bit RCA proposed by applying the FAFA2 obtained the best results in terms of the number of cycles and energy consumption compared to the exact 8-bit RCA [16], and the approximate 8-bit RCA contains the first approach’s approximate full adder. This improved approximate 8-bit RCA reduces the number of cycles and energy consumption by up to 44.642% and 51.589%, respectively.
The number of memristors and computational steps of the proposed circuits in the structures of $1^{st}$ and $2^{nd}$ scenarios are also compared with the IMPLY-based exact and approximate circuits [1, 21, 22] in Table 8. As mentioned earlier, it is impossible to directly compare the energy consumption of the proposed circuits with similar structures in [1, 21, 22] due to the difference in the memristor model applied. The number of memristors required in the $1^{st}$ and $2^{nd}$ implementation scenarios structures, based on the exact and approximate circuits proposed in [1, 21, 22], is 45.72% less than the FELIX-based exact and approximate structures. Suppose it is necessary to store the inputs in the work memristors using the IMPLY design method and reuse them later. In that case, the number of memristors required in the exact and approximate 8-bit structures ($1^{st}$ and $2^{nd}$ scenarios) will be equal to 35, which is equal to the number of memristors required in the FELIX-based exact and approximate 8-bit adders. The number of computational steps of the proposed circuits in the aforementioned dual scenarios is at most 82.39% and at least 64.66% less than the exact and approximate 8-bit structures based on IMPLY-based approximate full adders [1, 21, 22]. The improvement percentages reported in Table 8 do not consider the computational steps of copying input data into the surplus memristors in the IMPLY method.
@@ 4.2 Error-Analysis Results
Table 9: Error analysis results of the proposed approximate 8-bit RCA in two different scenarios.
@@
In the proposed approximate full adder (FAFA), the $C_{out}$ is calculated precisely, which means that the error does not propagate from the least significant approximate full adders to more significant approximate and exact full adders. Accordingly, and based on the numbers reported in Table 9, the computational accuracy of the proposed circuit in both scenarios is higher than the computational accuracy of the 8-bit approximate arithmetic structures based on the IMPLY-based approximate full adders [1, 22] in $1^{st}$ and $2^{nd}$ scenarios.
@@ 4.3 Simulations Of Various Image Processing Applications
Approximate computing can be applied in some applications that depend on human senses because human vision and hearing abilities are limited [41]. Approximate adders can be employed in error-resilient applications such as image processing. In this paper, the FAFA has been evaluated in three different image processing applications. These three applications are image addition, grayscale filter, and motion detection.
@@
Figure 8: The results of image addition application: (a) Cameraman, (b) Rice, (c) Exact addition, (d) Approximate addition in the $1^{st}$ scenario, and (e) Approximate addition in the $2^{nd}$ scenario.
@@
Figure 9: Motion detection application results: (a) First image, (b) Second image, (c) Exact result, (d) Approximate motion detection in the $1^{st}$ scenario, and (e) Approximate motion detection in the $2^{nd}$ scenario.
@@
Figure 10: The results of applying the grayscale filter: (a) Original RGB image, (b) exact result, (c) approximate grayscale filter in the $1^{st}$ scenario, and (d) approximate grayscale filter in the $2^{nd}$ scenario.
@@
According to the numbers reported in Tables 10 and 12, the quality of output images in image addition and grayscale filter applications resulting from the arithmetic structures constructed of the proposed approximate full adder is superior to the other circuits [1, 22] in both scenarios, indicating that the computational accuracy has decreased to an acceptable level. In the motion detection application, FAFA ranked second after SAFAN in both scenarios, surpassing the circuits proposed in [22] regarding image quality evaluation criteria. The high quality of output images indicates that the computational accuracy has decreased to an acceptable range, and the proposed circuit can be effectively employed in error-resilient computational applications.
@@ 5 Conclusion
This paper introduces a serial FELIX-based approximate full adder cell for IMC of error-resilient applications in two different implementation approaches. According to the advantages of approximate computing, the reduction of the number of memristors of the FAFA1 and FAFA2 compared to the exact FELIX-based full adder is 14.28% and 28.57%, respectively, and the energy consumption reduction is 73.735% and 81.754%, respectively. Also, the number of cycles in both approaches has decreased by 66.66% compared to the exact design. In this paper, MED, NMED, PSNR, SSIM, and MSSIM criteria were used to evaluate the accuracy of the FAFA in two different scenarios and three different image processing applications. In the $1^{st}$ and $2^{nd}$ scenarios, the exact full adder is applied in three MSBs; the FAFA adder is applied in five and four LSBs. The evaluations and simulations of the FAFA show that, in all applications and scenarios, PSNR is greater than 30 dB, and acceptable results were obtained regarding the accuracy criteria. In the mentioned scenarios, by applying the FAFA2, the number of cycles and energy consumption are improved by a maximum of 44.642% and 51.589%, respectively.

### v3

@@ 4.1 Circuit-Level Simulation And Analysis
* Not Calculated.
As shown in Table 8, the memristors’ count in both approximate scenarios and implementation approaches equals the memristors’ count in the exact 8-bit structure [16]. In both scenarios, the approximate 8-bit RCA proposed by applying the FAFA2 obtained the best results in terms of the number of cycles and energy consumption compared to the exact 8-bit RCA [16], and the approximate 8-bit RCA contains the first approach’s approximate full adder. This improved approximate 8-bit RCA reduces the number of cycles and energy consumption by up to 45.32% and 51.096%, respectively.
The number of memristors and computational steps of the proposed circuits in the structures of $1^{st}$ and $2^{nd}$ scenarios are also compared with the IMPLY-based exact and approximate circuits [1, 21, 22, 32, 33, 37, 38] in Table 8. As mentioned earlier, it is impossible to directly compare the energy consumption of the proposed circuits with similar structures in [1, 21, 22, 32, 33, 37, 38] due to the difference in the memristor model applied. The number of memristors required in the $1^{st}$ and $2^{nd}$ implementation scenarios structures, based on the exact and approximate circuits proposed in [1, 21, 22, 32, 33, 37, 38], is 14.29%-32.14% less than the FELIX-based exact and approximate structures. Suppose it is necessary to store the inputs in the work memristors using the IMPLY design method and reuse them later. In that case, the number of memristors required in the exact and approximate 8-bit structures ($1^{st}$ and $2^{nd}$ scenarios) will be equal to 36, which is 8 memristors more than the number of memristors required in the FELIX-based exact and approximate 8-bit adders. The number of computational steps of the proposed circuits in the aforementioned dual scenarios is at most 80.11% and at least 46.05% less than the exact and approximate 8-bit structures based on IMPLY-based approximate full adders [1, 21, 22, 32, 33, 37, 38]. The improvement percentages reported in Table 8 do not consider the computational steps of copying input data into the surplus memristors in the IMPLY method.
@@ 4.2 Error-Analysis Results
In the proposed approximate full adder (FAFA), the $C_{out}$ is calculated precisely, which means that the error does not propagate from the least significant approximate full adders to more significant approximate and exact full adders. Accordingly, and based on the numbers reported in Table 9, the computational accuracy of the proposed circuit in both scenarios is higher than the computational accuracy of the 8-bit approximate arithmetic structures based on the IMPLY-based approximate full adders proposed in [1, 22, 38] and the SINC design in [37] in $1^{st}$ and $2^{nd}$ scenarios.
Table 9: Error analysis results of the approximate 8-bit RCAs in two different scenarios.
@@ 4.3 Simulations Of Various Image Processing Applications
Approximate computing can be applied in some applications that depend on human senses because human vision and hearing abilities are limited [45]. Approximate adders can be employed in error-resilient applications such as image processing. In this paper, the FAFA has been evaluated in four different image processing applications. These four applications are image addition, grayscale filter, motion detection, and average pooling.
@@
With the expansion of neural network use across various processing applications, Convolutional Neural Networks (CNNs) are among the most popular neural network types in image processing [47]. One of the effective computational operations in CNNs is pooling, which reduces the input dimensions. Using the pooling function increases computational efficiency while preserving the main features of the input [47]. One pooling technique used in CNNs is average pooling, which reduces the effect of input noise [47]. To implement this operation, a $\times2$ kernel is sliding on the input image and the average value of every 4 pixels is stored as the final result. If the input of this process is an $n\times n$ image, the output is an $\frac{n}{2}\times\frac{n}{2}$ image with the features of the original image preserved. The proposed approximate full adders are used in the approximate 8-bit RCAs of the $1^{st}$ and $2^{nd}$ scenarios to compute the final sum of the average pooling function.
The simulation results of these applications are written in Tables 10-13. Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index Measure (SSIM), and Mean SSIM (MSSIM) criteria have been used to assess the quality of output images resulting from these applications. More details about calculating these criteria are explained in [7, 48]. It is mentioned in [49] that when the PSNR of the output images is greater than 30 dB, the quality of the output images is acceptable. As seen in Tables 10-13, the FAFA in both scenarios and all four applications has achieved PSNR greater than 30 dB, indicating the acceptable quality of the output images. Simulation outputs with SSIM values greater than 0.9 are acceptable for image processing applications [50]. Our simulation results in both scenarios achieve SSIM values greater than 0.9 in all cases except the second scenario in the motion detection application. Also, MSSIM values in all applications and scenarios are greater than 0.9, which is acceptable for output images. The simulation results of these four applications applying FAFA in both scenarios are shown in Figures 7-10.
Figure 7: The results of image addition application: (a) cameraman, (b) rice, (c) exact addition, (d) approximate addition in the $1^{st}$ scenario, and (e) approximate addition in the $2^{nd}$ scenario (quantitative metrics are listed in Table 10).
Figure 8: Motion detection application results: (a) first image, (b) second image, (c) exact result, (d) approximate motion detection in the $1^{st}$ scenario, and (e) approximate motion detection in the $2^{nd}$ scenario (quantitative metrics are listed in Table 11).
Figure 9: The results of applying the grayscale filter: (a) reference RGB image, (b) exact result, (c) approximate grayscale filter in the $1^{st}$ scenario, and (d) approximate grayscale filter in the $2^{nd}$ scenario (quantitative metrics are listed in Table 12).
Figure 10: Average pooling application results: (a) reference grayscale image, (b) exact result, (c) approximate average pooling in the $1^{st}$ scenario, and (e) approximate average pooling in the $2^{nd}$ scenario (quantitative metrics are listed in Table 13).
@@
Table 13: Image quality evaluation criteria in average pooling application.
8-bit Serial Approximate RCA PSNR (dB) SSIM MSSIM
SIAFA1 [22] 34.7106 0.983 0.9824
SIAFA2 [22] 35.8146 0.9684 0.9673
SIAFA3 [22] 41.6697 0.9886 0.9882
SIAFA4 [22] 35.0342 0.9787 0.9772
SAFAN [1] 35.1589 0.9852 0.985
SAID1 [38] 39.6548 0.9802 0.979
SAID2 [38] 40.5835 0.986 0.9858
SINC [37] 34.7844 0.9835 0.9824
SINC+ [37] 40.5606 0.984 0.9833
FAFA1 and FAFA2 40.9997 0.9844 0.9838
SIAFA1 [22] 33.5434 0.9469 0.9435
SIAFA2 [22] 29.2172 0.8884 0.8825
SIAFA3 [22] 34.8578 0.9602 0.9575
SIAFA4 [22] 27.5895 0.9302 0.9236
SAFAN [1] 28.9039 0.9526 0.95
SAID1 [38] 32.9688 0.9397 0.9346
SAID2 [38] 34.7106 0.9625 0.9612
SINC [37] 27.2967 0.9453 0.9401
SINC+ [37] 33.4426 0.955 0.9522
FAFA1 and FAFA2 34.2609 0.9404 0.9374
Figure 11: Image quality degradation (PSNR) versus error analysis metrics (NMED), the trade-off illustration between approximation error and perceptual quality: (a) $1^{st}$ scenario, and (b) $2^{nd}$ scenario.
According to the numbers reported in Tables 10-13, the quality of output images in grayscale filter applications resulting from the arithmetic structures constructed of the proposed approximate full adder is superior to the other circuits [1, 22, 37, 38] in both scenarios, indicating that the computational accuracy has decreased to an acceptable level. In the image addition, motion detection, and average pooling applications, FAFA ranked among the top three in both scenarios, and its image quality met the criteria for visually acceptable images. The high quality of the output images indicates that the computational accuracy has decreased to an acceptable level, and the proposed circuit can be effectively employed in error-resilient computational applications.
In Figure 11, two diagrams are drawn to examine the connection between the error metrics (NMED) and the image quality metrics of FAFA and state-of-the art [1, 22, 37, 38], which are applied in the architectures of scenarios 1 and 2. The average PSNR of four applications was applied to draw this diagram. The results show that there is an inverse trend between these two evaluation criteria; the higher the NMED, the lower the image quality metric. By examining the diagrams in Figure 11, it can be concluded that the proposed full adder cell and SINC+ [37] are the best options for use in the error-resilient applications.
Considering the diagrams in Figure 11 and their purposes, it can be concluded that the approximate RCAs (in scenarios 1 and 2) implemented using SINC+ [37] and FAFA have the highest computational accuracy. It is worth noting that the main goal of designing the FAFA was to improve the circuit evaluation criteria while maintaining acceptable computational accuracy, which has been achieved, as reported.
@@ 5 Conclusion
This paper introduces a serial FELIX-based approximate full adder cell for IMC of error-resilient applications in two different implementation approaches. According to the advantages of approximate computing, the reduction of the number of memristors of the FAFA1 and FAFA2 compared to the exact FELIX-based full adder is 14.28% and 28.57%, respectively, and the energy consumption reduction is 73.735% and 81.754%, respectively. Also, the number of cycles in both approaches has decreased by 66.66% compared to the exact design. In this paper, MED, NMED, PSNR, SSIM, and MSSIM criteria were used to evaluate the accuracy of the FAFA in two different scenarios and three different image processing and average pooling applications. In the $1^{st}$ and $2^{nd}$ scenarios, the exact full adder is applied in four and three MSBs; the FAFA adder is applied in four and five LSBs. The evaluations and simulations of the FAFA show that, in all applications and scenarios, PSNR is greater than 30 dB, and acceptable results were obtained regarding the accuracy criteria. In the mentioned scenarios, by applying the FAFA2, the number of cycles and energy consumption are improved by a maximum of 45.32% and 51.096%, respectively.
Appendix A: Application-level simulation-supplementary simulations and analyzes
In subsection 4.3, the functionality of the proposed full adder cells (FAFA1 and FAFA2) was analyzed across various image processing applications, including average pooling, which is applicable in machine learning and CNNs. The applicability of the previous works [1, 22, 37, 38] was also examined in these applications. The generated outputs were evaluated using various criteria, and the results indicated that the proposed circuits can be efficiently applied to computational structures, such as RCAs, in scenarios 1 and 2.
To gain greater confidence in the performance of the proposed circuits in the aforementioned applications, the circuits were incorporated into 8-bit approximate adder structures in both scenarios, and their performance in each application was evaluated on a set of standard images from two datasets [46, 51]. The images in the dataset [51] were used for image addition, grayscale filter, and average pooling, and three categories of images in the dataset [46] were used for the motion detection application. Image quality evaluation criteria were calculated for each of the images in all four applications, and the average results obtained are reported separately in Tables A1-A4 for each of the scenarios. By comparing the mentioned average results of the image quality metrics with those in Tables 10-13, it can be concluded that the proposed circuits, when applied to the 8-bit approximate RCAs in scenarios 1 and 2, provide acceptable functionality across all four applications and produce outputs of acceptable quality. Therefore, considering the results reported in Tables 10-13 and A1-A4, it can be concluded that the use of the proposed approximate circuits, if used deliberately in larger arithmetic structures, improves the circuit evaluation criteria, such as energy consumption, and produces outputs with acceptable accuracy and appropriate quality. The ultimate goal of using approximate computing and designing approximate arithmetic circuits is also in accordance with this procedure and the results obtained.
Table A1: Average Image quality metrics of the image addition application supplementary simulation based on the pictures of [51].
8-bit Serial Approximate RCA PSNR (dB) SSIM MSSIM
FAFA1 and FAFA2 39.445 0.969 0.97
FAFA1 and FAFA2 33.695 0.912 0.908
Table A2: Average Image quality metrics of the motion detection application supplementary simulation based on the pictures of [46].
8-bit Serial Approximate RCA PSNR (dB) SSIM MSSIM
FAFA1 and FAFA2 39.712 0.948 0.957
FAFA1 and FAFA2 34.177 0.891 0.903
Table A3: Average Image quality metrics of the grayscale filter application supplementary simulation based on the pictures of [51].
8-bit Serial Approximate RCA PSNR (dB) SSIM MSSIM
FAFA1 and FAFA2 41.998 0.976 0.99
FAFA1 and FAFA2 36.034 0.93 0.97
Table A4: Average Image quality metrics of the average pooling application supplementary simulation based on the pictures of [51].
8-bit Serial Approximate RCA PSNR (dB) SSIM MSSIM
FAFA1 and FAFA2 40.407 0.973 0.974
FAFA1 and FAFA2 33.858 0.912 0.915
Statements & Declarations
Funding
No funds, grants, or other support was received.
Competing interests
The authors declare no competing interests.
Author Contributions
Seyed Erfan Fatemieh: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Writing - Original Draft, Writing – Review & Editing, Visualization.
Samane Asgari: Conceptualization, Methodology, Software, Validation, Formal Analysis, Investigation, Writing - Original Draft, Writing - Review & Editing, Visualization.
Mohammad Reza Reshadinezhad: Validation, Formal Analysis, Investigation, Resources, Writing - Review & Editing, Supervision, Project Administration.
Data Availability Statement
All data have been generated through the methods described in this paper.